# LLM as GNN: Graph Vocabulary Learning for Text-Attributed Graph Foundation Models

### v1

@@ 4.2 Graph Inference Module
The graph inference module seeks to unify diverse graphs and tasks and acquire transferable knowledge through multi-instruction fine-tuning using LLMs. Due to distinct vocabularies, existing methods, which represent nodes as OOV tokens, suffer from semantic misalignment between ID-based and language-based embeddings in task-oriented prompts, constraining their transferability. To tackle this limitation, we introduce a novel language-based graph vocabulary that bridges this incompatibility, enabling massive readable and coherent instructions to support LLM inference.
Graph Vocabulary Learning. Learning a transferable graph vocabulary, whose fundamental units can represent each node, is key to building GFMs. Its effectiveness depends on three essential criteria: expressiveness, transferability, and scalability. Since each node has been associated with a textual representation that encapsulates its core semantics and local structure, we intuitively propose a universal language-based graph vocabulary using these rich representations. In this vocabulary, each node is represented by a finite sequence of language tokens, i.e. a language-based ID, defined as follows:
@@
where each graph node $v_{i}$ is assigned a sequence as ${T}_{i}=\left(t_{i}^{1},t_{i}^{2},\ldots,t_{i}^{m_{i}}\right)$, where $t_{i}^{k}\in\mathcal{X}$ and $\mathcal{X}$ is a dictionary of general natural language tokens. Along this line, all nodes in $V$ have language-based IDs as $T=\left\{T_{0},T_{1},\dots,T_{|V|-1}\right\}$. Consequently, natural language tokens form the fundamental building blocks of our graph vocabulary, with structured token sequences representing graph nodes, analogous to words in standard lexicons.
Our graph vocabulary meets all expected criteria: (a) Expressiveness. Language-based IDs preserve rich semantic and structural information from an open-world setting. (b) Transferability. Like human lexicons, it shares natural language as a common foundation, ensuring inherent cross-graph transferability. (c) Scalability. Any node, whether previously seen or not, can dynamically generate its language-based ID, ensuring compatibility with existing nodes and mitigating token explosion.
Instruction Fine-Tuning. We employ a multi-instruction fine-tuning framework to incorporate various graphs and tasks (Chung et al., 2024; Wei et al., 2022). As illustrated in Figure 4, we index nodes from the graph vocabulary and embed their language-based IDs into task-oriented prompt templates to construct completed instructions $\mathcal{T}$:
@@
which include language-based IDs of the central nodes and their respective local structure $T$ from $G$, depending on specific graph-centric tasks (e.g., node classification or link prediction). Thus, these instructions are fully readable and coherent, as they are composed entirely of natural language tokens. Afterward, we convert all the question-answering tasks into a unified text-to-text format for LLM fine-tuning (Mishra et al., 2022). Given target sequences $Y$, the loss function is computed as:
@@
where $\text{Pr}\left(Y_{j}\mid\mathcal{T},Y_{<j}\right)$ is the probability of the $j$-th token $Y_{j}$ in the output sequence $Y$, conditioned on the instruction $\mathcal{T}$ and all previous tokens $Y_{<j}=(Y_{1},Y_{2},\dots,Y_{j-1})$. Using T5 (Raffel et al., 2020), FLAN  (Wei et al., 2022), or Llama  (Touvron et al., 2023) as the backbone, we can co-train a unified LLM across diverse graphs and tasks by fine-tuning, which enables the acquisition of open-world global knowledge and inclusive accommodation of unseen graphs or tasks.
Constrained Decoding via Prefix Tree Search. To mitigate LLM hallucination in link prediction, we introduce a constrained decoding method using a prefix tree search strategy to regulate LLM outputs  (Cao et al., 2021; Tan et al., 2024). Specifically, we craft a prefix tree from language-based IDs of all candidate nodes, where each tree node represents a natural language token. Each unique path from the root to a leaf corresponds to a specific language-based ID. During autoregressive generation, token output is restricted to a valid path within the prefix tree, ensuring predictions align with actual graph nodes and eliminating hallucinations. This effectiveness is attributed to the discrete nature of language-based IDs, further highlighting the flexibility of the proposed graph vocabulary.
@@
PromptGFM (Llama3) 92.42 85.32 94.65 83.78 86.72 86.61 84.66
5 Experiments
@@
$\bullet$ RQ1: How does PromptGFM perform on supervised node classification and link prediction?
$\bullet$ RQ2: How is the transferability across diverse graphs and tasks as a versatile GFM?
$\bullet$ RQ3: How does each module contribute to overall performance of PromptGFM?
$\bullet$ RQ4: What affect the GNN replication?
Datasets. We present the following datasets covering three domains: (a) computer science: Cora (McCallum et al., 2000), Citeseer (Giles et al., 1998), Ogbn-arxiv (Hu et al., 2020), and WikiCS  (Mernyei and Cangea, 2020); (b) e-commerce: Photo and History (Ni et al., 2019; Yan et al., 2023); (c) biomedical: PubMed (Sen et al., 2008). Details can be found in Appendix A.
Baselines. We make comprehensive comparisons with existing methods in four categories: (1) Graph-agnostic methods. We consider the MLP model without graph structure. (2) GNN-based methods. We employ four fundamental GNN models: GCN  (Kipf and Welling, 2017), GAT  (Velickovic et al., 2017), GraphSAGE  (Hamilton et al., 2017), and ReVGNN  (Li et al., 2021). We also explore SGFormer  (Wu et al., 2023) and NodeFormer  (Wu et al., 2022), which leverage transformer architectures to model graph data. (3) BERT-based methods. We utilize BERT (Devlin et al., 2019) and Sentence-BERT  (Reimers and Gurevych, 2019) to obtain textual representations for downstream tasks. (4) GNN-LLM Integration methods. Following the proposed taxonomy, we select GraphPrompter  (Liu et al., 2024b) as an instance of using GNNs to enhance LLMs. OFA  (Liu et al., 2024a) and ENGINE  (Zhu et al., 2024) are examples of leveraging LLMs for GNNs. Besides, we incorporate LLaGA  (Chen et al., 2024a) as an attempt of implementing LLM as GNN. We provide baseline details in Appendix B.
Reproduction Settings. We implement PromptGFM in PyTorch and run all experiments on four NVIDIA RTX A6000 GPUs. The graph understanding module leverages OpenAI’s GPT-4o mini while we fine-tune a Flan-T5 or Llama3-8B in the graph inference module. For evaluation metrics, we use accuracy and Macro-F1 (abbreviated as M-F1) for node classification, and accuracy and HR@1 for link prediction, respectively. We provide further implementation details in Appendix E.
5.1 Performance Comparison (RQ1)
Main Results. We train each model independently from scratch on a single graph and compare their performance. Tables 1 and 2 present the results for node classification and link prediction accuracy, respectively. PromptGFM achieves substantial improvements over state-of-the-art models, leading to the following insights. First, graph-based models outperform graph-agnostic ones, highlighting the importance of structural information. Second, our method surpasses OFA, ENGINE, and GraphPrompter in node classification, partly due to their decoupled architectures in both GNN-for-LLM and LLM-for-GNN paradigms. Third, PromptGFM outperforms LLaGA, which relies on templates to understand graph semantics and structures. The results suggest that heuristic prompts alone fail to capture sufficient high-order signals without an actual GNN mechanism. In contrast, PromptGFM demonstrates the potential of leveraging LLMs as GNNs through a prompt-based GNN, establishing a novel paradigm for graph-text alignment in TAGs.
Generative Link Prediction. Following a transductive setting, we partition the graph by links, create an input graph from the training set, and finally predict unseen connections for test nodes, with existing nodes as potential candidates. From Table 3, PromptGFM consistently outperforms traditional GNN models. Unfortunately, existing GNN-LLM approaches overlook this setting due to their reliance on OOV token embeddings, preventing LLM outputs from mapping to specific nodes and leading to unresolved hallucination issues. In contrast, our framework represents nodes as finite token sequences, thereby allowing constrained decoding to regulate LLM outputs. This further underscores the critical role of our graph vocabulary and the flexibility and scalability of PromptGFM.
5.2 Cross-graph Transferability (RQ2)
We examine cross-graph transferability in both intra-domain and inter-domain scenarios. Following  (Chen et al., 2024b), we explore two settings: (1) Pre-training, where models transfer to unseen datasets without prior exposure during training, and (2) Co-training, where available target data is also integrated for joint training with auxiliary datasets. This results in four distinct transfer paradigms.
Intra-domain Cross-graph Transferability. As shown in Table 4, all settings surpass the off-the-shelf LLM. Meanwhile, incorporating Arxiv yields greater gains than Cora, suggesting that larger and richer datasets enhance transferability more effectively. In the pre-training setting, significant improvements over direct inference with an off-the-shelf LLM highlights strong zero-shot transferability within the computer science domain, while similar results on co-training further confirm effective knowledge transfer. These findings underscore the potential to collect more graph data and build more comprehensive and knowledgeable GFMs.
@@ 5.2 Cross-graph Transferability (RQ2)
Inter-domain Cross-graph Transferability. We outline cross-domain results in Table 6. In the pre-training setting, while incorporating a single external graph enhances performance, adding multiple graphs can lead to negative transfer. More critically, in the co-training setting under supervised learning, performance deteriorates as more cross-domain data is introduced. This decline stems from domain-specific knowledge conflicts, potentially caused by catastrophic forgetting or incompatible hyperparameters that fail to accommodate to varying domain distributions. Although PromptGFM provides a unified interface for diverse graphs and tasks, we leave robust positive knowledge transfer as future work, further discussed in Section 7.
[…]
Table 4: Intra-domain cross-graph generalization. We transfer from the computer science domain to the Citeseer dataset for node classification task. The first row indicates direct inference using off-the-shelf LLM.
@@
[…]
Setting Cora Citeseer PubMed
zero-shot 18.54 12.16 27.64 17.10 39.12 39.84
[…]
ST-NC 91.72 90.06 84.49 80.13 90.67 91.82
5.3 Ablation Studies (RQ3)
To assess the contribution of each module, we introduce three ablated variants: (1) w/o understanding, which removes the prompt-based GNN; (2) w/o inference, which eliminates multi-instruction fine-tuning; and (3) w/o both, which directly prompts LLMs using raw textual attributes. Figure 5 reports node classification accuracy, where our full model consistently outperforms all variants. First, the drop in w/o understanding indicates that omitting GNN replication leads to a loss of semantic and structural information. Second, w/o inference shows a significant decline, emphasizing the crucial role of LLM fine-tuning in integrating domain-specific knowledge. Lastly, w/o both yields the worst results, underscoring the synergy between the understanding and inference modules.
Table 6: Inter-domain cross-graph generalization. We transfer from the computer science domain to biomedical domain (i.e. PubMed) on node classification.
@@ 5.3 Ablation Studies (RQ3)
5.4 Exploration Studies (RQ4).
We analyze factors affecting prompt-based GNNs. Since GNNs are permutation-invariant, we evaluate LLM-powered prompt-based GNNs using Flan-T5 with two variants: (1) shuffle nodes, which randomizes neighbor order, and (2) shuffle tokens, which injects position-specific tokens to nodes and shuffle them instead. As illustrated in Table 7, regardless of the shuffle method, node classification accuracy exhibits only minor fluctuations, having negligible impact on the stability of prompt-based GNNs.
Additionally, we investigate the impact of layer depth in Appendix F. Furthermore, we present a case study on layer-by-layer representations in Appendix G.1, and compare language-based IDs with keywords from citation datasets in Appendix G.2.
@@
Table 7: Permutation sensitivity of prompt-based GNNs.
Variant Cora Citeseer
shuffle nodes 90.64 $\pm$1.12 83.79 $\pm$ 0.89
shuffle tokens 90.55 $\pm$1.03 84.28 $\pm$1.06

### v2

@@ 4.2 Graph Inference Module
The graph inference module seeks to unify diverse graphs and tasks and acquire transferable knowledge through multi-instruction fine-tuning with LLMs. Existing methods, which represent nodes as OOV tokens, suffer from semantic misalignment between ID-based and language-based embeddings in task-oriented prompts due to distinct vocabularies, constraining their transferability. To tackle this limitation, we introduce a novel language-based graph vocabulary that bridges this incompatibility, enabling the generation of large-scale, coherent, and readable instructions to support LLM inference.
Graph Vocabulary. Learning a transferable graph vocabulary, whose fundamental units can represent each node, is critical for building effective GFMs. Its utility hinges on three key criteria: expressiveness, transferability, and scalability. Since each node has been equipped with a textual representation in the graph understanding module that captures its core semantics and local structure, we intuitively construct a universal language-based graph vocabulary based on these concise yet informative representations. In this vocabulary, each node is represented by a finite sequence of language tokens, known as a language-based ID, defined as:
@@
where each graph node $v_{i}$ is assigned a sequence as ${T}_{i}=\left(t_{i}^{1},t_{i}^{2},\ldots,t_{i}^{m_{i}}\right)$, where $t_{i}^{k}\in\mathcal{X}$ and $\mathcal{X}$ is a dictionary of general natural language tokens. Collectively, the language-based IDs for all nodes in $V$ are denoted as $T=\left\{T_{0},T_{1},\dots,T_{|V|-1}\right\}$. Along this line, natural language tokens serve as the fundamental building blocks of our graph vocabulary, with structured token sequences representing each graph node, analogous to words in standard lexicons.
Our vocabulary satisfies all expected criteria: (a) Expressiveness. Language-based IDs preserve rich semantic and structural information from open-world settings. (b) Transferability. Like human lexicons, the vocabulary is grounded in natural language, providing a shared foundation that ensures cross-graph interpretability and transferability. (c) Scalability. Any node, whether previously seen or not, can dynamically generate its language-based ID, maintaining compatibility with existing nodes while preventing token explosion. We analyze vocabulary coverage and uniqueness in Section 5.3.
Multi-Instruction Fine-Tuning. As shown in Figure 4, we employ a multi-instruction fine-tuning framework to incorporate diverse graphs and tasks [9, 51]. We index nodes from the vocabulary and embed their language-based IDs into task-oriented templates to construct completed instructions $\mathcal{T}$:
@@
which include language-based IDs of the central nodes, along with its corresponding local structure from $G$ depending on specific graph-centric tasks (e.g., node classification or link prediction). As these instructions are composed entirely of natural language tokens, they are fully readable and coherent. All graph-centric tasks are then reformulated in a text-to-text format, allowing the LLM to be fine-tuned within a unified instruction learning paradigm [34]. Given a target sequence $Y$, the loss function can be computed as:
@@
where $\text{Pr}\left(Y_{j}\mid\mathcal{T},Y_{<j}\right)$ denotes the probability of the $j$-th token $Y_{j}$ in the output sequence $Y$, conditioned on the instruction $\mathcal{T}$ and all previous tokens $Y_{<j}=(Y_{1},Y_{2},\dots,Y_{j-1})$. To support broad generalization, we fine-tune LLM backbones such as T5 [39], FLAN [51], and Llama [45] to co-train a unified model across diverse graphs and tasks. As a result, the model acquires open-world knowledge and generalizes effectively to previously unseen graphs and tasks.
Figure 4: An instance of graph inference module in link prediction, where language-based IDs are indexed from the graph vocabulary to generate readable instructions with task-oriented templates. A multi-instruction fine-tuning framework is employed to unify graphs and tasks and learn transferable knowledge for GFMs.
Constrained Decoding via Prefix Tree Search. To mitigate LLM hallucination, we introduce a constrained decoding strategy based on a prefix tree search to regulate LLM outputs  [4, 43]. Taking generative link prediction as an example, we craft a prefix tree from language-based IDs of all candidate nodes, where each tree node corresponds to a natural language token and each root-to-leaf path uniquely identifies a valid node ID. During autoregressive generation, token selection is constrained to a valid path within the prefix tree, ensuring that predictions align with actual graph nodes and eliminating hallucinations. This effectiveness can be attributed to the discrete nature of language-based IDs, further highlighting the flexibility of our proposed graph vocabulary.
5 Experiments
$\bullet$ RQ1: How does PromptGFM perform on node classification and link prediction tasks?
$\bullet$ RQ2: How transferable is PromptGFM across diverse graphs and tasks as a general-purpose GFM?
$\bullet$ RQ3: How effective is the proposed graph vocabulary in terms of quantitative evaluation?
$\bullet$ RQ4: How does each module contribute to the overall performance of PromptGFM?
$\bullet$ RQ5: What is the computational efficiency in graph understanding and inference modules?
$\bullet$ RQ6: What factors influence the effectiveness of GNN replication within PromptGFM?
Datasets. We present the following datasets covering three domains: (a) computer science: Cora [32], Citeseer [13], Ogbn-arxiv [18], and WikiCS  [33]; (b) e-commerce: Photo and History [36, 57]; (c) biomedical: PubMed [42]. Details can be found in Appendix A.
Baselines. We make comprehensive comparisons with existing methods in four categories: (1) Graph-agnostic methods. We consider the MLP model without graph structure. (2) GNN-based methods. We employ four fundamental GNN models: GCN  [23], GAT [46], GraphSAGE [16], and ReVGNN [24]. We also explore SGFormer [54] and NodeFormer [53], which leverage Transformer architectures to model graph data. (3) LM-only methods. We utilize BERT [10], Sentence-BERT  [40], and Llama3-8B [45] to address downstream tasks. (4) GNN-LLM Integration methods. Following the proposed taxonomy, GraphPrompter  [29] is an instance of using GNNs to enhance LLMs. OFA  [27] and ENGINE  [62] are examples of leveraging LLMs for GNNs. Besides, we incorporate LLaGA  [6] as an attempt of implementing LLM as GNN. Details can be found in Appendix B.
@@ 5 Experiments
PromptGFM (Llama3-8B) 92.42 85.32 94.65 83.78 86.72 86.61 84.66
$p$-value 4e-3 2.73e-8 - 4.48e-10 4.85e-7 3.62e-2 9.93e-4
@@
Reproduction Settings. We conduct all experiments in PyTorch on four NVIDIA RTX A6000 GPUs. The graph understanding module uses OpenAI’s GPT-4o mini while we fine-tune a Flan-T5 or Llama3-8B in the graph inference module. We employ 10-fold cross-validation and compute average results with standard deviations. For evaluation, we use accuracy and Macro-F1 (M-F1) for node classification, and accuracy and HR@1 for link prediction. To align with the TAG setting, BERT is used for node initialization across all applicable baselines. More details are provided in Appendix E.
5.1 Performance Comparison (RQ1)
Main Results. We train each model independently from scratch on a single graph and compare their performance. Tables 1 and 3 summarize the results for node classification and link prediction accuracy, respectively. PromptGFM achieves substantial improvements over state-of-the-art models, leading to the following observations. (a) Graph-based models consistently outperform graph-agnostic ones, highlighting the importance of structural information. (b) Our method surpasses OFA, ENGINE, and GraphPrompter on node classification, suggesting the limitations of their decoupled architectures in both GNN-for-LLM and LLM-for-GNN paradigms. (c) PromptGFM also surpasses LLaGA, which relies on templates to implicitly understand graph semantics and structures. These results indicate that heuristic prompting alone is insufficient to capture high-order dependencies without a true GNN mechanism. In contrast, PromptGFM effectively replicates message passing through a prompt-based GNN, demonstrating the feasibility of using LLMs as GNNs and establishing a new paradigm for graph-text alignment.
Generative Link Prediction. Following a transductive setting, we partition the graph by links, construct an input graph from the training edges, and predict missing connections for test nodes, using existing nodes as candidate targets. From Table 3, PromptGFM consistently outperforms traditional GNN models in the generative link prediction. Unfortunately, existing GNN-LLM methods fail to address this task, as their reliance on OOV token embeddings prevents LLM outputs from being directly mapped to specific nodes, often resulting in hallucinated predictions. In contrast, our work represents nodes as finite token sequences, allowing constrained decoding via prefix tree search to regulate LLM outputs. This further highlights the essence of our graph vocabulary and demonstrates the flexibility of PromptGFM in generative scenarios.
Table 4: Intra-domain cross-graph transferability
to Citeseer within the computer science domain.
@@ 5.1 Performance Comparison (RQ1)
Table 5: Inter-domain cross-graph transferability from e-commerce to biomedical domain (i.e. PubMed).
Source Pre-training Co-training
History Photo Acc($\uparrow$) M-F1($\uparrow$) Acc($\uparrow$) M-F1($\uparrow$)
39.12 39.84 90.67 91.82
46.24 47.41 88.52 89.39
44.21 44.93 85.34 85.57
45.18 45.65 86.22 87.03
5.2 Cross-graph and Cross-task Transferability (RQ2)
We investigate cross-graph transferability in both intra-domain and inter-domain scenarios. Following  [7], we explore two settings: (1) Pre-training, where models are evaluated on entirely unseen datasets without prior exposure during training, and (2) Co-training, where a portion of the target dataset is available and jointly used with auxiliary datasets for training. These two dimensions yield four distinct transfer paradigms, enabling a comprehensive assessment of cross-graph transferability.
Intra-domain Cross-graph Transferability. As illustrated in Table 5, all settings surpass direct inference using an off-the-shelf LLM. Notably, incorporating Arxiv yields greater gains than Cora, suggesting that larger and richer datasets enhance transferability more effectively. In the pre-training setting, substantial improvements over direct inference demonstrate strong zero-shot transferability within the domain. A similar, albeit smaller, trend is observed in the co-training setting, further validating the effectiveness of intra-domain knowledge transfer. These findings underscore the potential to collect more graph data and build more knowledgeable and transferable GFMs.
Table 6: Inter-domain cross-graph transferability from computer science to biomedical domain (i.e. PubMed).
@@ 5.2 Cross-graph and Cross-task Transferability (RQ2)
Inter-domain Cross-graph Transferability. We report cross-domain transfer results from e-commerce and computer science to the biomedical domain in Tables 5 and 6, respectively. In the pre-training setting, while incorporating a single external graph enhances performance, adding multiple source graphs may lead to negative transfer. More critically, in the co-training setting under supervised learning, performance consistently deteriorates as more cross-domain data is introduced. This decline stems from domain-specific knowledge conflicts, potentially caused by catastrophic forgetting or incompatible hyperparameters that fail to adapt to varying domain distributions [7, 59]. While PromptGFM offers a unified interface for diverse graphs and tasks, achieving robust and reliable cross-domain transfer remains a challenge, which we leave for future exploration.
[…]
Setting Cora Citeseer PubMed
zero-shot 18.54 12.16 27.64 17.10 39.12 39.84
[…]
ST-NC 91.72 90.06 84.49 80.13 90.67 91.82
[…]
5.3 Exploration Studies
Vocabulary Analysis (RQ3). We evaluate the quality of our graph vocabulary using two metrics in Table 8. (a) Coverage. We compute the proportion of nodes that can be matched to IDs. Since the vocabulary is language-based, any node’s semantics can, in principle, be expressed through a combination of language tokens. Thus, every node is successfully assigned a valid ID, confirming complete coverage. (b) Uniqueness. We measure the entropy of token distributions across nodes in both individual and merged graphs. Higher entropy indicates more diverse and fine-grained token usage, indicating stronger uniqueness and representational capacity. Notably, the merged vocabulary (“ALL”) achieves the highest entropy (12.2761), showing that integrating multiple graphs improves vocabulary expressiveness and discriminability, which reinforces the necessity of building GFMs.
Ablation Studies (RQ4). We construct three ablated variants to assess the contribution of each component: (a) w/o understanding, which removes the prompt-based GNN; (b) w/o inference, which eliminates multi-instruction fine-tuning; and (c) w/o both, which directly prompts LLMs using raw textual attributes. Figure 6 reports the node classification accuracy, where our full model consistently outperforms all variants. Specifically, the performance drop in w/o understanding highlights the importance of replicating GNN workflow to preserve both semantic and structural signals. The sharp decline in w/o inference further emphasizes the necessity of LLM fine-tuning for incorporating domain-specific knowledge. Finally, w/o both yields the worst results, underscoring the complementarity between the graph understanding and inference modules.
@@ 5.3 Exploration Studies
Figure 6: Impact of layer numbers.
Table 8: Quantitative graph vocabulary analysis on coverage and entropy.
Metric Cora Citeseer PubMed Ogbn-arxiv History Photo WikiCS ALL
Coverage (%) 100.00 100.00 100.00 100.00 100.00 100.00 100.00 100.00
Entropy 9.1116 9.4561 8.4643 10.7464 10.3752 11.3272 9.3825 12.2761
Table 9: Efficiency analysis in terms of API cost (USD), number of tokens (M), training time (min), and power consumption (kWh).
Source Understanding Inference
Cora Citeseer Photo API Cost # Tokens Time Power
0.188 1.25 16.47 0.07
0.189 1.29 21.41 0.08
1.213 8.08 83.67 0.34
N/A N/A 35.31 0.16
N/A N/A 117.56 1.04
Efficiency Analysis (RQ5). We evaluate the computational and monetary cost of PromptGFM. As summarized in Table 9, while PromptGFM incurs an initial API cost during the graph understanding phase, this overhead is one-time and remains manageable with modern infrastructure. Specifically, each graph is processed only once to generate language-based node representations, which can be cached and reused throughout the inference pipeline. As a result, when additional graphs are introduced for training, there is no need to repeatedly call the LLM API to regenerate node representations. In contrast, traditional end-to-end models often require retraining from scratch when the graph structure changes, leading to higher cumulative cost. These characteristics make the computational overhead acceptable and amortizable, especially in dynamic or large-scale graph scenarios, where repeated retraining would otherwise be prohibitively expensive.
Prompt-based GNN Analysis (RQ6). We examine how the number of layers in our prompt-based GNN affects performance. As shown in Figure 6, increasing the number of layers initially improves results by capturing broader context and higher-order dependencies. However, beyond a certain depth, performance plateaus or declines due to over-smoothing, a trend consistent with traditional GNNs. Optimal performance is achieved with 3 layers on Cora and 2 layers on Citeseer. This suggests that textual representations can be propagated over the graph similarly to numerical embeddings, effectively capturing both semantics and structure. We further provide a case study on layer-wise representations in Appendix H.1, and compare language-based IDs with citation keywords in Appendix H.2. Additionally, we investigate permutation sensitivity of the prompts in Appendix G.